OpenStudy (anonymous):

A medical test has a 95% accuracy of detecting a Condition Z if the person has it. It also has a 97% chance to indicate that the person does not have the condition if they really don't have it. If the incidence rate of this disease is 10 out of every 100: What is the probability that a person chosen at random will both test positive and actually have the disease (i.e., get a true positive)?

OpenStudy (anonymous):

Events are going to be A - tests positive (false positive or true positive) B - has desease

OpenStudy (anonymous):

And the question asks for \[ \Pr(A\cap B) \]

OpenStudy (anonymous):

"95% accuracy of detecting a Condition Z if the person has it" This means \[ \Pr(A|B) = 0.95 \]

OpenStudy (anonymous):

"incidence rate of this disease is 10 out of every 100"\[ \Pr(B) = \frac{10}{100} \]

OpenStudy (anonymous):

"It also has a 97% chance to indicate that the person does not have the condition if they really don't have it"\[ \Pr(A^C|B^C) = 0.97 \]

OpenStudy (anonymous):

Actually, I'm a bit unsure about parts of this, because it seems t hey game too much information.

OpenStudy (anonymous):

Well there was another question: What is the probability that a person chosen at random will test positive but not have the disease (i.e., get a false positive)? But I left it out because I figured once I figured out how to do it I'd be able to figure the 2nd one

OpenStudy (anonymous):

Oh, ok.

OpenStudy (anonymous):

Do you know this formula: \[ \Pr(A|B) = \frac{\Pr(A\cap B)}{\Pr(B)} \]

OpenStudy (anonymous):

We can use it to solve for \(\Pr(A\cap B)\).

OpenStudy (anonymous):

I only know how to do intersections when there's multiple numbers

OpenStudy (anonymous):

Sets I mean

OpenStudy (anonymous):

Do you understand the formula I gave?

OpenStudy (anonymous):

Yeah, but I'm not sure how to do the intersection when it's only 2 numbers

OpenStudy (anonymous):

Well \[ \Pr(A\cap B) \]means probability of event A and B. \[ \Pr(A|B) \]means probability of A if we know B is true

OpenStudy (anonymous):

\[ \Pr(A|B)=\frac{\Pr(A\cap B)}{\Pr(B)} \]We can multiply by \(\Pr(B)\) and then: \[ \Pr(A\cap B) = \Pr(A|B)\times \Pr(B) \]

OpenStudy (anonymous):

But how do I do the upside down U part?

OpenStudy (anonymous):

Are you talking about \[ \Pr(A\cup B) = \Pr(A) +\Pr(B) - \Pr(A\cap B) \]

OpenStudy (anonymous):

The Pr(A *Flipped U* B)

OpenStudy (anonymous):

Oh, you do `\(\cap\)` \(\cap\)

OpenStudy (anonymous):

Well, how do I do that part of the equation though

OpenStudy (anonymous):

The intersection part

OpenStudy (anonymous):

I think it's 9.5%?

OpenStudy (anonymous):

Yeah, that looks right.

OpenStudy (anonymous):

Thanks :)